# CP-LLM: Context and Pixel Aware Large Language Model for Video Quality Assessment

### v1

@@ 1 Introduction
Figure 1: Videos with intentional effects but incorrectly identified as quality defects by traditional VQA models (CoINVQ wang2021rich , FastVQA wu2022fast , and DOVER wu2022disentangling ), while MLLM’s assessment aligns better with human perception. The first two videos contain intentional low-light and background blurring. The third video presents an astronomy video containing complex visual textures. The last video shows a pixel-art style game characterized by inherent blockiness. All quality scores are calibrated to $[1,5]$, where $1$, $2$, $3$, $4$, $5$ correspond to bad, poor, fair, good, and excellent respectively.
@@ 3.1.1 High-level Encoder
We denote a video sequence as $\bm{x}=\{\bm{x}_{i}\}_{i=0}^{N-1}$, where $\bm{x}_{i}\in\mathbb{R}^{H\times W\times 3}$ represents the $i$-th frame, $H$ and $W$ are the original frame height and width, and $N$ is the total number of frames in the sequence. We first uniformly sample a sparse set of $M$ key frames, denoted as $\bm{y}=\{\bm{y}_{i}\}_{i=0}^{M-1}$, from the original sequence $\bm{x}$. Each selected key frame $\bm{y}_{i}\in\mathbb{R}^{H\times W\times 3}$ undergoes preprocessing. A vision encoder $\bm{f}_{h}(\cdot)$ processes the input frame $\bm{y}_{i}$ after bilinear resizing to $H_{h}\times W_{h}$. This encoder outputs a sequence of contextualized visual embeddings $\bm{v}_{i}=\bm{f}_{h}(\bm{y}_{i})$. $\bm{u}_{i}$ is then are then projected through a linear layer to align with a target dimension $D$ (specifically, the embedding dimension of language decoder), resulting $\bm{u}_{i}\in\mathbb{R}^{T_{h}\times D}$.
@@ 3.1.2 Low-level Encoder
The frame $\bm{v}_{i}$ is resized using bilinear interpolation to a fixed resolution $H_{l}\times W_{l}$. Subsequently, we extract $K$ non-overlapping patches, $\mathcal{P}_{i}=\{\bm{p}_{i,j}\}_{j=0}^{K-1}$, uniformly from the resized frame, where each patch $\bm{p}_{i,j}\in\mathbb{R}^{H_{p}\times W_{p}\times 3}$ has a size $H_{p}\times W_{p}$. These $K$ patches for each key frame $\bm{y}_{i}$ are then fed into a vision Transformer (ViT) model dosovitskiy2020image , denoted as $\bm{f}_{l}(\cdot)$. $\bm{f}_{l}$ processes the patch sequence $\mathcal{P}_{i}$ and outputs a corresponding sequence of feature embeddings $\bm{h}_{i}=\{\bm{h}_{i,j}\}_{j=0}^{K-1}$, where $\bm{h}_{i,j}\in\mathbb{R}^{T_{q}\times D_{q}}$, $T_{q}$ is the number of tokens and $D_{q}$ is the feature dimension of $\bm{f}_{l}$. The features $\bm{h}_{i,j}$ are then projected through a linear layer to align with a target dimension $D$ as well. To obtain a fixed-size representation for the frame while reducing the sequence length, we apply average pooling across the patch dimension and followed by a layer normalization, resulting in the final low-level embeddings $\bm{v}_{i}\in\mathbb{R}^{T_{l}\times D}$.
@@ 3.2 Robust Multi-task Training Pipeline
where $\mathcal{B}$ denotes the batch size, $q_{i}$ represents the ground-truth quality score (either a mean opinion score (MOS) for an original video or a pseudo-MOS for a generated variant), and $\hat{q}_{i}$ is the predicted score. Superscripts $(1)$ and $(2)$ indicate the first and second video in the $i$-th pair, respectively.
@@ 4.1 Experimental Setups
Datasets. The proposed model, CP-LLM, was trained on LSVQ ying2021patch and augmented with compression variants. These variants were generated by encoding the original LSVQ videos using the H.264 codec at $20$ distinct constant rate factor (CRF) levels, ranging from $15$ to $53$. A foundational assumption in this process is that the perceptual quality of a variant encoded at a higher CRF level (i.e., lower bitrate) does not exceed that of a variant encoded at a lower CRF level (i.e., higher bitrate). To train the language decoder component of CP-LLM, videos from LSVQ were annotated using Gemini 2.0, prompted with: “Based on the provided video frames, create a three-sentence summary. First, describe the visual content of the video in detail. Second, identify the style of the video. Third, assess the technical quality.” These generated annotations constitute a crucial part of the training data for the language decoder. For evaluation, the performance of CP-LLM was assessed on five diverse UGC datasets selected for their varied content types and distortion characteristics: YT-Compress wang2019youtube , YouTubeUGC wang2019youtube , LIVE-YT-Gaming yu2022subjective , Shorts-SDR wang2024youtube , and Shorts-HDR2SDR wang2024youtube .
@@
Implementation Details. We set the number of key frames $M=5$ per video for LSVQ, Shorts-SDR and Shorts-HDR2SDR, $M=8$ for LIVE-YT-Gaming, $M=19$ for YT-Compress and YouTube-UGC. Key frames were processed via two pipelines: 1) bilinear resizing to $448\times 448$ pixels ignoring aspect ratio for the SigLIP chen2023pali encoder, and 2) bilinear resizing preserving aspect ratio to fit within $540\times 1080$ pixels, followed by cropping $K=8$ non-overlapping $224\times 224$ pixel patches for the low-level vision encoder VideoPrism zhao2024videoprism . The maximum sequence length was set to $L=512$ tokens. We fine-tuned the Gemma team2024gemma language decoder using low-rank adaptation (LoRA) hu2022lora , applying a rank of $R=4$ to its query, key, and value attention projections. This fine-tuning process utilized the same prompt that guided the initial data annotation phase. During this process, the high-level vision encoder remained fixed, while the entire low-level vision encoder and the quality regression head underwent full fine-tuning. Optimization was performed with the multi-task loss with equal weights using the Adam optimizer kingma2014adam for $500$ epochs with a batch size of $128$ and an initial learning rate of $\eta=1\times 10^{-4}$ on a $16$-core TPU v5.
@@ 4.2 Experiment Results
Table 1: MOS/DMOS correlations and flip rates evaluated on the YT-Compress dataset. The top-$2$ results are highlighted in bold and underline.
Table 1 compares the VQA model accuracy and robustness on the YT-Compress dataset, evaluating by SRCC/PLCC on mean opinion scores (MOS) and differential mean opinion scores (DMOS), and flip rate (FR) on pairs of compression variants. We can see CP-LLM achieves superior performance on all three dimensions. Specifically, it surpasses Q-Align, another MLLM-based approach. This suggests that CP-LLM not only excels in making accurate quality assessments for single videos but is also sensitive at detecting distortions within the same contents.
The cross-dataset MOS correlations are shown in Table 2. These four datasets represent highly challenging and recent VQA benchmarks, with content that diverges substantially from the LSVQ training set. CP-LLM achieves a top-$2$ ranking across these four datasets, demonstrating satisfactory generalization capabilities.
Table 3 shows the flip rate results observed across $20$ distinct compression variants of the LSVQ-test-1080p dataset, aiming to evaluate the fine-grained distortion sensitivity. CP-LLM showcases superior sensitivity to pixel-level distortions, outperforming all other models. This highlights the efficacy of CP-LLM’s architecture and comprehensive training strategy in capturing pixel-level artifacts. Notably, prominent models such as Gemini 2.0, considered a leading MLLM, and DOVER, a SOTA VQA model, exhibit limitations in reliably distinguishing small pixel-level distortions. This suggests that in the absence of the proposed quality-related pairwise training, even highly capable models may struggle with such fine-grained perceptual distinctions.
@@
Table 2: Cross-dataset testing of our model against four competing models, all retrained on the official training split of the large-scale LSVQ dataset and tested on other VQA datasets without fine-tuning. The top-$2$ results on each dataset are highlighted in bold and underline.
@@
Table 3: Flip rate test results for our proposed model versus five competing models, evaluated on LSVQ-test-1080p and its compression variants. “Answer / Score” denotes if the model’s comparison output is a textual answer or a numerical score; “Diff” indicates the CRF difference between compression pairs. The top-$2$ results are highlighted in bold and underline.

### v2

@@ 1 Introduction
Figure 1: Videos with intentional effects but incorrectly identified as quality defects by traditional VQA models (CoINVQ wang2021rich , FastVQA wu2022fast , and DOVER wu2022disentangling ), while MLLM’s assessment aligns better with human perception. The first two videos contain intentional low-light and background blurring. The third video presents an astronomy video containing complex visual textures. The last video shows a pixel-art style game characterized by inherent blockiness. All quality scores are calibrated to [1,5][1,5][ 1 , 5 ], where 111, 222, 333, 444, 555 correspond to bad, poor, fair, good, and excellent respectively.
@@ 3.1.1 High-level Encoder
We denote a video sequence as $\bm{x}=\{\bm{x}_{i}\}_{i=0}^{N-1}$, where $\bm{x}_{i}\in\mathbb{R}^{H\times W\times 3}$ represents the $i$-th frame, $H$ and $W$ are the original frame height and width, and $N$ is the total number of frames in the sequence. We first uniformly sample a sparse set of $M$ key frames, denoted as $\bm{y}=\{\bm{y}_{i}\}_{i=0}^{M-1}$, from the original sequence $\bm{x}$. Each selected key frame $\bm{y}_{i}\in\mathbb{R}^{H\times W\times 3}$ undergoes preprocessing. A vision encoder $\bm{f}_{h}(\cdot)$ processes the input frame $\bm{y}_{i}$ after bilinear resizing to $H_{h}\times W_{h}$. This encoder outputs a sequence of contextualized visual embeddings $\bm{u}_{i}=\bm{f}_{h}(\bm{y}_{i})$. $\bm{u}_{i}$ is then are then projected through a linear layer to align with a target dimension $D$ (specifically, the embedding dimension of language decoder), resulting $\bm{u}_{i}\in\mathbb{R}^{T_{h}\times D}$.
@@ 3.1.2 Low-level Encoder
The frame $\bm{y}_{i}$ is resized using bilinear interpolation to a fixed resolution $H_{l}\times W_{l}$. Subsequently, we extract $K$ non-overlapping patches, $\mathcal{P}_{i}=\{\bm{p}_{i,j}\}_{j=0}^{K-1}$, uniformly from the resized frame, where each patch $\bm{p}_{i,j}\in\mathbb{R}^{H_{p}\times W_{p}\times 3}$ has a size $H_{p}\times W_{p}$. These $K$ patches for each key frame $\bm{y}_{i}$ are then fed into a vision Transformer (ViT) model dosovitskiy2020image , denoted as $\bm{f}_{l}(\cdot)$. $\bm{f}_{l}$ processes the patch sequence $\mathcal{P}_{i}$ and outputs a corresponding sequence of feature embeddings $\bm{h}_{i}=\{\bm{h}_{i,j}\}_{j=0}^{K-1}$, where $\bm{h}_{i,j}\in\mathbb{R}^{T_{q}\times D_{q}}$, $T_{q}$ is the number of tokens and $D_{q}$ is the feature dimension of $\bm{f}_{l}$. The features $\bm{h}_{i,j}$ are then projected through a linear layer to align with a target dimension $D$ as well. To obtain a fixed-size representation for the frame while reducing the sequence length, we apply average pooling across the patch dimension and followed by a layer normalization, resulting in the final low-level embeddings $\bm{v}_{i}\in\mathbb{R}^{T_{l}\times D}$.
@@ 3.2 Robust Multi-task Training Pipeline
where $\mathcal{B}$ denotes the batch size, $q_{i}$ represents the ground-truth quality score (either a mean opinion score (MOS) for an original video or a pseudo-MOS for a generated variant), and $\hat{q}_{i}$ is the predicted score. Superscripts (1)(1)( 1 ) and (2)(2)( 2 ) indicate the first and second video in the $i$-th pair, respectively.
@@ 4.1 Experimental Setups
Datasets. The proposed model, CP-LLM, was trained on LSVQ ying2021patch and augmented with compression variants. These variants were generated by encoding the original LSVQ videos using the H.264 codec at 202020 distinct constant rate factor (CRF) levels, ranging from 151515 to 535353. A foundational assumption in this process is that the perceptual quality of a variant encoded at a higher CRF level (i.e., lower bitrate) does not exceed that of a variant encoded at a lower CRF level (i.e., higher bitrate). To train the language decoder component of CP-LLM, videos from LSVQ were annotated using Gemini 2.0, prompted with: “Based on the provided video frames, create a three-sentence summary. First, describe the visual content of the video in detail. Second, identify the style of the video. Third, assess the technical quality.” These generated annotations constitute a crucial part of the training data for the language decoder. For evaluation, the performance of CP-LLM was assessed on five diverse UGC datasets selected for their varied content types and distortion characteristics: YT-Compress wang2019youtube , YouTubeUGC wang2019youtube , LIVE-YT-Gaming yu2022subjective , Shorts-SDR wang2024youtube , and Shorts-HDR2SDR wang2024youtube .
@@
Implementation Details. We set the number of key frames $M=5$ per video for LSVQ, Shorts-SDR and Shorts-HDR2SDR, $M=8$ for LIVE-YT-Gaming, $M=19$ for YT-Compress and YouTube-UGC. Key frames were processed via two pipelines: 1) bilinear resizing to $\times 448$ 448 pixels ignoring aspect ratio for the SigLIP chen2023pali encoder, and 2) bilinear resizing preserving aspect ratio to fit within 540×1080540$\times$ 1080540 × 1080 pixels, followed by cropping $K=8$ non-overlapping $\times 224$ 224 pixel patches for the low-level vision encoder VideoPrism zhao2024videoprism . The maximum sequence length was set to $L=512$ tokens. We fine-tuned the Gemma team2024gemma language decoder using low-rank adaptation (LoRA) hu2022lora , applying a rank of $R=4$ to its query, key, and value attention projections. This fine-tuning process utilized the same prompt that guided the initial data annotation phase. During this process, the high-level vision encoder remained fixed, while the entire low-level vision encoder and the quality regression head underwent full fine-tuning. Optimization was performed with the multi-task loss with equal weights using the Adam optimizer kingma2014adam for 500500500 epochs with a batch size of 128128128 and an initial learning rate of $\eta=1\times 10^{-4}$ on a 161616-core TPU v5.
@@ 4.2 Experiment Results
Table 1: MOS/DMOS correlations and flip rates evaluated on the YT-Compress dataset. The top-222 results are highlighted in bold and underline.
Table 1 compares the VQA model accuracy and robustness on the YT-Compress dataset, evaluating by SRCC/PLCC on mean opinion scores (MOS) and differential mean opinion scores (DMOS), and flip rate (FR) on pairs of compression variants. We can see CP-LLM achieves superior performance on all three dimensions. Specifically, it surpasses Q-Align, another MLLM-based approach. This suggests that CP-LLM not only excels in making accurate quality assessments for single videos but is also sensitive at detecting distortions within the same contents. Moreover, although trained exclusively with H.264 compression, the model demonstrates notable robustness to the VP9 compression present in the YT-Compress dataset.
The cross-dataset MOS correlations are shown in Table 2. These four datasets represent highly challenging and recent VQA benchmarks, with content that diverges substantially from the LSVQ training set. CP-LLM achieves a top-222 ranking across these four datasets, demonstrating satisfactory generalization capabilities.
Table 3 shows the flip rate results observed across 202020 distinct compression variants of the LSVQ-test-1080p dataset, aiming to evaluate the fine-grained distortion sensitivity. CP-LLM showcases superior sensitivity to pixel-level distortions, outperforming all other models. This highlights the efficacy of CP-LLM’s architecture and comprehensive training strategy in capturing pixel-level artifacts. Notably, prominent models such as Gemini 2.0, considered a leading MLLM, and DOVER, a SOTA VQA model, exhibit limitations in reliably distinguishing small pixel-level distortions. This suggests that in the absence of the proposed quality-related pairwise training, even highly capable models may struggle with such fine-grained perceptual distinctions.
@@
Table 2: Cross-dataset testing of our model against four competing models, all retrained on the official training split of the large-scale LSVQ dataset and tested on other VQA datasets without fine-tuning. The top-222 results on each dataset are highlighted in bold and underline.
@@
Table 3: Flip rate test results for our proposed model versus five competing models, evaluated on LSVQ-test-1080p and its compression variants. “Answer / Score” denotes if the model’s comparison output is a textual answer or a numerical score; “Diff” indicates the CRF difference between compression pairs. The top-222 results are highlighted in bold and underline.